OpenStudy (anonymous):

Maximize Q = xy2, where x and y are positive numbers, such that x + y2 = 4.

OpenStudy (anonymous):

does y2 means y squared?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If so then the solution of the problem is: You want to maximize Q, that means you need to find the derivative and critical points of Q. The problem here is that Q is in terms of x and y. You should make it in term of only one of them. So we are going to use the second equation to do so. \[x+y^2=4 \implies y^2=4-x\] substitute for the value of y^2 into Q, you get: \[Q=xy^2=x(4-x)=4x-x^2\]

OpenStudy (anonymous):

Do you know how to maximize 4x-x^2?

OpenStudy (anonymous):

I plug test points into 4x-x^2?

OpenStudy (anonymous):

the critical number is 4?

OpenStudy (anonymous):

^^ yeah exactly.

OpenStudy (anonymous):

so the max is 3?

OpenStudy (anonymous):

you solved for x when the derivative is zero, and you got tgat x=4. this is your only critical point, so it's either maximum or minimum. Which one do you think it is?

OpenStudy (anonymous):

oooh. i dont know

OpenStudy (anonymous):

hey wait!!

OpenStudy (anonymous):

I made a mistake :(

OpenStudy (anonymous):

lol oh no

OpenStudy (anonymous):

the critical point is not x=4

OpenStudy (anonymous):

I didn't make it, you did and I didn't see it :P

OpenStudy (anonymous):

Don't worry.. It's so easy. Do you know how to find the derivative of 4x-x^2?

OpenStudy (anonymous):

4- 2x

OpenStudy (anonymous):

Right!! Now solve for x when 4-2x=0.

OpenStudy (anonymous):

2!

OpenStudy (anonymous):

haha so hard

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Now 2 is your critical point, and at x=2 Q is a maximum value.

OpenStudy (anonymous):

So now just plug x=2 in Q=4x-x^2, and you're done!!

OpenStudy (anonymous):

yay thanks

OpenStudy (anonymous):

What is the maximum value?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

:) I hope you can do any similar problem next time!

OpenStudy (anonymous):

thanks for your help!!

OpenStudy (anonymous):

no problem.